# No Vulnerability Data, No Problem: Towards Predicting Mean Time To Remediate In Open Source Software Dependencies

### v2

@@ 9. Conclusion and Future Works
In this study, we introduced two dependency update metrics, MTTUdep and MTTRdep, to quantify the updatedness of dependencies in open-source software packages. Our large-scale empirical analysis across the npm, PyPI, and Cargo ecosystems demonstrated that MTTUdep can serve as a proxy for MTTRdep when vulnerability information is unavailable. Furthermore, our statistical analysis highlighted the relationships between package characteristics and dependency update behavior, providing actionable insights for developers, maintainers, and software supply chain researchers. Future research can explore additional factors influencing dependency update practices, such as the severity of vulnerabilities, organizational policies, or developer incentives. Expanding this analysis to other ecosystems, combining with transitive dependencies, and incorporating qualitative insights from developers could further refine our understanding of dependency updatedness.

### v3

@@ 9. Conclusion and Future Works
In this study, we introduced two dependency update metrics, MTTUdep and MTTRdep, to quantify the updatedness of dependencies in open-source software packages. Our large-scale empirical analysis across the npm, PyPI, and Cargo ecosystems demonstrated that MTTUdep can serve partially as a proxy for MTTRdep when vulnerability information is unavailable. Furthermore, our statistical analysis highlighted the relationships between package characteristics and dependency update behavior, providing actionable insights for developers, maintainers, and software supply chain researchers. Future research can explore additional factors influencing dependency update practices, such as the severity of vulnerabilities, organizational policies, or developer incentives. Expanding this analysis to other ecosystems, combining with transitive dependencies, and incorporating qualitative insights from developers could further refine our understanding of dependency updatedness.
Data Availability
The code and data for the analysis in this paper are all available in our replication package in Zenodo (zenodo-artifact, ). It is currently restricted for reviewers only, but we will make it public upon acceptance.